OpenStudy (anonymous):

I have a distance problem and I know the equation is d=rt. I'm just not sure how to solve the problem: A flight crew flew 420 km in 3 h with a tailwind. Flying against the wind, the flight crew flew 440 km in 4 h. Find the rate of the flight crew in calm air and the rate of the wind.

OpenStudy (anonymous):

the total distance traveled divide by total amount time should give you an average of how much traveled per hour

OpenStudy (anonymous):

so it would be 420/3? and 440/4 to get the rate?

OpenStudy (anonymous):

since you are looking for a rate of travel in calm air, i think (420+440)/(3+4) should do it

OpenStudy (anonymous):

ok Thank you so much!!

OpenStudy (anonymous):

How would I find the rate of the wind?

OpenStudy (anonymous):

is the rate of the wind considered same for both cases?

OpenStudy (anonymous):

I guess it would be....the answer I get for the first part is 122.8 km/h. Do I use that answer to determine the rate of the wind?

OpenStudy (anonymous):

one way to find out is - since you have the rate of travel at calm wind, you can subtract rate of travel at either tail wind or against the wind from calm wind

OpenStudy (anonymous):

wind=15 km/hr, aircraft speed = 125 km/hr

OpenStudy (anonymous):

i don't think you can just add the rate of travel since the time at which air plane travels is different

OpenStudy (anonymous):

i think the rate of wind might be different for both cases

OpenStudy (anonymous):

@robtobey, how did you arrive at 125 km/h = aircraft speed and 15 km/hr = wind? Was the original equation wrong or was mymath wrong? :)

OpenStudy (anonymous):

msayer3, There is no indication from the problem statement that the wind velocity changed from my point of view.

OpenStudy (anonymous):

yeah, that's what i'm wondering.

OpenStudy (anonymous):

I will post the solution in a few minutes.

OpenStudy (anonymous):

ok...thank yo very much

OpenStudy (anonymous):

r*t = d r = d/t Let x = the wind speed First flight leg: \[r+x==\frac{d}{t}\text{/.} \{d\to 420,t\to 3\} \]r + x = 140 Second flight leg:\[r-x=\frac{d}{t}\text{/.} \{d\to 440,t\to 4\} \]r - x = 110 2 r = 140 + 110 r = 125 r + x =140 125 + x = 140 x = 15

OpenStudy (anonymous):

I use Mathematica for the calculations /. means sort or, "replace the symbols as follows"

OpenStudy (anonymous):

Ok...I think I got it now. Thank you again for your help.

OpenStudy (anonymous):

Glad I could help you out.